Two groups of people went into the shop. The first group bought four hotdogs and six juices and paid 71 dollars. The second group bought three hotdogs and seven juices and paid 67 dollars. What were the prices for the hotdogs and juices?
Answer by jorel555(1290)

Let h be hot dogs, j be juices. Then:
4h+6j=71
3h+7j=67
So;
12h+18j=213
12h+28j=268
10j=268-213=55
j=$5.50
h=$9.50 ☺☺☺☺